OpenStudy (anonymous):

Evaluate the expression: sin 30 degrees cos 60 degrees + sin 60 degrees cos 30 degrees

OpenStudy (cherylim23):

Recall the double angle formula here, thus expression = sin (30 + 60) = sin90 = 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what about (cos 30 degrees)^2 - (sin30degrees)^2

OpenStudy (cherylim23):

Likewise double angle, cos(30+30)= cos(60) = 1/2

OpenStudy (anonymous):

hmm my teacher never said anything about this double angle formula!!! ugh summer school. Thanks

OpenStudy (lgbasallote):

this isnt double angle actually..

OpenStudy (lgbasallote):

are you familiar with this @Lanik \[\large \sin (\alpha + \beta) = \sin \alpha \cos \beta + \sin \beta \cos \alpha\]

OpenStudy (cherylim23):

Oops typo my bad. The first case should have been compound angle formula.

OpenStudy (lgbasallote):

yep better

OpenStudy (lgbasallote):

i didnt notice the question changed lol

OpenStudy (lgbasallote):

@Lanik if you're interested to "derive" the double angle formula.. \[\cos (\alpha + \beta) = \cos \alpha \cos \beta - \sin \alpha \sin \beta\] \[\text{If} \quad \alpha = \beta\] \[\cos (\alpha + \alpha) = \cos \alpha \cos \alpha - \sin \alpha \sin alpha\] \[\cos(2\alpha) = \cos^2 \alpha - \sin^2 \alpha\]